An automatic filling machine is used to fill 1-liter bottles of cola. The machine’s output is approximately normal with a mean of .97 liter and a standard deviation of .02 liter. Output is monitored using means of samples of 25 observations. Use Table-A.


a.
Determine upper and lower control limits that will include roughly 97 percent of the sample means when the process is in control. (Do not round intermediate calculations. Round your answers to 4 decimal places.)



Upper control limits:
liter
Lower control limits:
liter

Answer by stanbon(75887)

An automatic filling machine is used to fill 1-liter bottles of cola.
The machine’s output is approximately normal with a mean of .97 liter and a standard deviation of .02 liter.
Output is monitored using means of samples of 25 observations. Use Table-A.
---------------------
a.
Determine upper and lower control limits that will include roughly 97 percent of the sample means when the process is in control. (Do not round intermediate calculations. Round your answers to 4 decimal places.)
-----
Find the z-values that have tails of 3%.
+-invNorm(0.03) = +-1.8808
--------
Upper control limits = 1.8808*.02+0.97 = 1.0076 liters
------------------
Lower control limits = -1.8808*0.02+0.97 = 0.9324 liter
------------
Cheers,
Stan H.
--------------